How do you solve aequations such as 4(e+4)=-12? Here is a list of problems I am having trouble with:
8-9y=35
8-12x=32
0=25x+75
3n-9=9
x/3-8=-12
So if you could just tell me how to solve this type of problem it would be greatly appreciated. Thanks

Answer by fractalier(6550)

In general, you want to isolate the variable...that is, you want to do whatever you can to get the letter by itself on one side of the equal sign and a number on the other side. To do that, you have to UNDO what is being done to the variable (again, in general). The cardinal rule in algebra is that whatever you do to one side of an equation, you MUST do to the other side. 100% of the time -- no exceptions.
So let's do this one:
8 - 9y = 35
Here I want to get y by itself. First I will subtract 8 from BOTH sides.
-9y = 27
Then I will undo the multiplication by -9, by dividing bth sides by -9...
-9y/-9 = 27/-9
y = -3
and so on...for
4(e + 4) = -12
we'll divide by 4 first
e + 4 = -3
then subtract 4
e = -7